OpenStudy (anonymous):

What is the equation of the line that passes through the point (2, 5) and has a slope of 2?

OpenStudy (the_fizicx99):

Put it in: Y - y1=M(x-x1)

OpenStudy (leoray):

Use the equation and plug the y's and x's.... y+y^1= (slope which is 2) (x-x^1)

OpenStudy (the_fizicx99):

What?!? ^